A student fires a cannonball vertically upwards. The cannonball returns to the ground after a 6.60s flight. Determine all unknowns and answer the following questions. Neglect drag and the initial height and horizontal motion of the cannonball. Use regular metric units (ie. meters).
1
Expert's answer
2020-12-10T11:04:04-0500
"v=0.5gT=0.5(9.8)(6.6)=32.3\frac{m}{s}\\
h=0.5(0.25)gT^2=0.5(0.25)(9.8)(6.6)^2=53.4\ m"